Bandar Industries Berhad of Malaysia manufactures sporting equipment. One of the company’s products, a football helmet for the North American market, requires a special plastic. During the quarter ending June 30, the company manufactured 3,800 helmets, using 2,546 kilograms of plastic. The plastic cost the company $16,804.

According to the standard cost card, each helmet should require 0.60 kilograms of plastic, at a cost of $7.00 per kilogram.

Required:

1. What is the standard quantity of kilograms of plastic (SQ) that is allowed to make 3,800 helmets?

2. What is the standard materials cost allowed (SQ × SP) to make 3,800 helmets?

3. What is the materials spending variance?

4. What is the materials price variance and the materials quantity variance?
Answer #1

1. Standard quantity = 3,800 x 0.60 = 2,280 kilograms

2. Standard material cost = 2,280 x 7 = $15,960

3. Material spending variance = standard cost - actual cost

= 15,960 - 16,804

= -844

4.

Material price variance = (SP - AP) ×AQ

= (7 - 6.60) × 2,546

= 1,018.4

Material quantity variance = (SQ - AQ) × SP

= (2,280 - 2,546) × 6.60

= -1,755.6